# Vision-driven Automated Mobile GUI Testing via Multimodal Large Language Model

### v1

@@ III-B Function-aware Explorer
III-B1 Function-based Testing History Construction
@@
The TEXT view depicts the higher-level catalog of testing progress, organized according to the functions. It includes a list of tested functions along with their visit counts, which can assist the MLLM in avoiding duplicate exploration and uncovering untouched functions. To accomplish this, during the testing process, the MLLM is prompted to infer what function is currently being tested through the “prompt of function inquiry”, as demonstrated in Fig 5 (b). This enables us to gather statistics on the tested functions
The IMAGE view offers a detailed snapshot of the testing records through screenshots, providing a granular view of the testing process. Specifically, it includes a list of detailed GUI pages, with each page marked with a widget (or a group of widgets), i.e., denoting the corresponding actions when exploring that GUI page, which is demonstrated in Fig. 5 (a). Additionally, we provide derived functions for each screenshot to further ensure alignment between visual and textual information. We also associate a unique identifier (ID) with the function name, representing different pages/steps within the same function.
Figure 6: Example of explorer prompt
III-B2 Explorer Prompt Generation
With the static GUI information (in Section III-A1) and the dynamic testing progress (in Section III-B1), we devise prompts for inputting into the MLLM.
Prompts related to GUI text and visual information are demonstrated in Fig. 6 with “GUI text info” and “legend of current GUI screenshots”, offering the MLLM an overview of the current GUI page.
Prompts related to action & feedback are illustrated in Fig. 6 with “querying general action”, “querying text input”, and “testing feedback”. We prompt the MLLM to identify what operation is required, and provide the output template to assist the MLLM in generating a desired executive command for app testing. It’s worth noting that we separate the patterns for querying general action (e.g., click) and text input (e.g., input certain text) respectively, to enable the generated commands to better match the widgets in the GUI page. For the feedback question, if we detect the widget ID does not correspond to its textual description, which is determined with the help of the alignment information in Section III-A, we inform the MLLM that there might be an error in its decision-making process and prompt it to retry. ‘widget number’ is not accurate, better to use ‘widget id’, and revise them in other places and the figures.
Prompts related to function-based history are depicted in Fig. 6 with “explored functions” and “legend of recently tested screenshots”. They are fetched from the testing history, where the former is directly obtained from the text view, and the latter is extracted from the image view, displaying the latest four tested GUI screenshots to provide the most relevant clues.
Prompts related to function inquiry are illustrated in Fig. 6 with “querying function”, prompting the MLLM to identify what function is currently tested, and the prompt also provides an output template to assist the MLLM in generating the function name and its status.
Figure 7: Example of sequence segmentation and bugs detected by VisionDroid
III-C Logic-aware Bug Detector
During the exploration process, intra-page non-crash bugs are directly detected using MLLM, triggered by the “bug detect” prompt shown in Fig. 6. This section will focus on describing the bug detector for inter-page non-crash bugs.
Once the exploration phase is complete, we leverage the entire exploration history for bug detection. However, the overall history is often too lengthy, intricate, and noisy for directly querying the MLLM. The direct manner is to organize the entire history into individual parts based on the functions inferred by MLLM during the exploration process. Yet the MLLM tends to yield relatively granular function names, often comprising only 2 to 3 operational steps. However, many functional bugs are typically associated with the business logic formed by such consecutive functions. For instance, as depicted in Fig. 7, a bug involves functions like “edit expense”, “delete a budget”, “check budget”, etc. To tackle this, we propose a logic-driven testing sequence segmentation method to partition the entire exploration history into logically cohesive parts.
III-C1 Logic-driven Testing Sequence Segmentation Algorithm
The principle of test sequence segmentation is that GUI pages with similar operational logic are considered as a sub-sequence, while GUI pages with different logic are considered as different sub-sequences. Based on this, we begin by constructing a graph based on the entire testing history. Each node represents an explored GUI page (i.e., a screenshot from the image view in the testing history as shown in Fig. 5). Each edge signifies the transition between nodes. We assign the edge weights based on the semantic similarity of their associated function names (i.e., the text below each screenshot as shown in Fig. 5 (b)). Specifically, we segment the textual function name into tokens, and calculate the cosine similarity between them. In this way, the edge weight between pairs of pages with completely different functions should be 0, e.g., Check budget-2 and Setting-1 in Fig. 7.
Based on the weighted directed graph, the logic-driven testing sequence segmentation algorithm is conducted with the Louvain algorithm [23] and optimized based on modularity $Q$ for sub-sequence partitioning.
$Q=\frac{1}{2m}\sum_{ij}\left[A_{ij}-\frac{1}{2m}\right]\delta(c_{i},c_{j})$
$A_{ij}$ represents the weight between node $i$ and node $j$. $m$ is the sum of the weights of the graph. $c_{i}$ and $c_{j}$ are the sub-sequence number for node $i$ and node $j$. $\delta$ is an indicator function, when $c_{i}=c_{j}$ the value is 1, otherwise, it is 0, used to determine whether two nodes belong to the same sub-sequence.
Figure 8: Example of bug detection prompt
III-C2 Detector Prompt Generation
With the segmented results, we proceed to feed each sub-sequence into the MLLM for bug detection. Specifically, we provide MLLM with the related screenshots in the form of IMAGE view as shown in Fig. 5.
Prompts related to testing sub-sequence are demonstrated in Fig. 8 with “APP info” and “legend of GUI screenshots with function name”, offering the MLLM with app’s overview and the details about the derived testing sub-sequence.
Prompts related to bug detection question are demonstrated in Fig. 8. We enable the MLLM to comprehend the semantics of the test sequence based on the provided information and assess the correctness of the test path, further inquiring if there are any bugs in the test sequence and identifying which step contains bugs.
III-D Enriching Detector Prompt with Examples
@@
It is usually difficult for MLLM to perform as well on domain-specific tasks as ours, and common practice would be employing the in-context learning schema to boost the performance. It provides the MLLM with examples to demonstrate what the instruction is, which enables the MLLM to better understand the task. Following the schema, along with the prompt for the detector, as described in Section III-C, we additionally provide the MLLM with examples of the non-crash logical bugs.
@@
To achieve this, we first build a basic example dataset of non-crash bugs from the issue reports of open-source mobile apps and continuously enlarge it. For each data instance, it records the bug description, bug screenshot and natural language described bug reproduction path which facilitates the MLLM understanding of what the non-crash bug is. We select 200 bugs (100 intra-page, 100 inter-page) with the highest quality and complete information as the example data. We also enrich the example dataset with the newly emerged non-crash bugs which truly trigger bugs during VisionDroid runs on various apps.
@@
Although examples can provide intuitive guidance to the LLM in accomplishing a task, excessive examples might mislead the LLM and cause the performance to decline. Therefore, we design a retrieval-based example selection method to choose the most suitable examples (i.e., most similar to the tested GUI page) for MLLM. The similarity comparison is based on the activity name of the tested page. Similar to previous works [24], We use Word2Vec [25] to encode the context information of each activity into a sentence embedding, and calculate the cosine similarity between the activity and each data instance in the example dataset. We choose the top-K data instance with the highest similarity score, and set K as 5.
III-E Implementation
@@
We implement VisionDroid based on the gpt-4 API which is released on the OpenAI website [^2] https://beta.openai.com/docs/models/GPT-4. It obtains the view hierarchy file of the current GUI page through UIAutomator [21] to extract text information of the input widgets. VisionDroid is implemented as a fully automated GUI testing tool, which uses or extends the following tools: VirtualBox and the Python pyvbox [26] for running and Android Debug Bridge (ADB) [27] for interacting with the app under test.
IV Experiment Design
  • RQ1: (Coverage and bugs detection performance) How effective of VisionDroid in testing coverage and bug detection?
@@ IV Experiment Design
For RQ1, we first present some general views of VisionDroid for coverage of the explorer and bug detection of the detector. Then we compare VisionDroid with commonly-used and state-of-the-art baseline approaches.
@@
  • RQ2: (Ablation Study) What is the contribution of the (sub-) modules of VisionDroid for coverage and bug detection performance?
@@
  • RQ3: (Usefulness Evaluation) How does our proposed VisionDroid work in real-world situations?
For RQ3, we use VisionDroid to automatically explore the app, and detect unseen non-crash functional bugs, and issue the detected bugs to the development team.
IV-A Datasets
@@
Considering the small number of apps in the baseline dataset, we collect a second dataset (Github dataset) following similar procedures as the baseline dataset. In detail, we crawl the 30 most popular apps of each category from Google Play [30], and we keep the ones with at least one update after May. 2024, resulting in 218 apps in 10 Google Play categories. Then we manually review each issue report and app, and filter them according to the following criteria: (1) the app wouldn’t constantly crash on the emulator; (2) it can run all baselines; (3) UIAutomator [21] can obtain the view hierarchy file and its GUI screenshot; (4) the issue report involve non-crash bug and it can be manually reproduced for validation; (6) the app is not used in the motivational study or example dataset construction. Please note that we follow the name of the app to ensure that there is no overlap between the datasets. Finally, 43 apps with 247 non-crash bugs remain for further experiments.
To avoid potential data leakage issues caused by the use of MLLM and further analyze the bug detection capability of VisionDroid, we construct a third dataset (Injection dataset) through bug injection, which can ensure that MLLM has not learned about these bugs in the open-source repositories. Specifically, we obtain the source code of the second dataset apps from GitHub and manually verify whether the app could conduce the bug injection by modifying the code through repackaging. Among them, 27 apps can run normally after bug injection. We conduct the injection based on the bug description, bug repair code, and bug root cause obtained through the pilot study in Section II, and implement it following the existing bug injection approach [31, 32, 33]. For the 9 summarized bug categories in Section II, we inject 5 non-crash bugs per category, resulting in a total of 45 bugs in 27 apps.
@@
Intra-page bug detection method with automated GUI testing
OwlEyes+A 0.29 0.27 92 18 0.20 0.14 0.41 0.37 117 57 0.49 0.23 0.42 0.40 77 7 0.09 0.16
NightHawk+A 0.29 0.27 86 21 0.24 0.16 0.41 0.37 99 66 0.67 0.27 0.42 0.40 41 8 0.20 0.18
DiffDroid+A 0.29 0.27 87 20 0.23 0.15 0.41 0.37 82 47 0.57 0.19 0.42 0.40 27 3 0.11 0.07
dVermin+A 0.29 0.27 67 6 0.09 0.05 0.41 0.37 51 10 0.20 0.04 0.42 0.40 18 2 0.11 0.04
OwlEyes+F 0.39 0.37 107 17 0.16 0.13 0.44 0.40 119 57 0.48 0.23 0.43 0.39 79 8 0.10 0.18
NightHawk+F 0.39 0.37 100 20 0.20 0.15 0.44 0.40 98 64 0.65 0.26 0.43 0.39 39 9 0.23 0.20
DiffDroid+F 0.39 0.37 94 20 0.21 0.15 0.44 0.40 77 51 0.66 0.21 0.43 0.39 28 4 0.14 0.09
dVermin+F 0.39 0.37 41 8 0.20 0.06 0.44 0.40 44 13 0.30 0.05 0.43 0.39 11 2 0.18 0.04
OwlEyes+H 0.38 0.35 97 12 0.12 0.09 0.44 0.40 97 54 0.56 0.22 0.43 0.39 71 7 0.10 0.16
NightHawk+H 0.38 0.35 88 16 0.18 0.12 0.44 0.40 94 58 0.62 0.23 0.43 0.39 69 7 0.10 0.16
DiffDroid+H 0.38 0.35 75 17 0.23 0.13 0.44 0.40 78 32 0.41 0.13 0.43 0.39 31 4 0.13 0.09
dVermin+H 0.38 0.35 33 5 0.15 0.04 0.44 0.40 27 17 0.63 0.07 0.43 0.39 11 2 0.18 0.04
OwlEyes+G 0.47 0.43 99 17 0.17 0.13 0.49 0.44 104 59 0.57 0.24 0.47 0.41 77 8 0.10 0.18
NightHawk+G 0.47 0.43 95 18 0.19 0.14 0.49 0.44 110 69 0.63 0.28 0.47 0.41 48 9 0.19 0.20
DiffDroid+G 0.47 0.43 85 16 0.19 0.12 0.49 0.44 88 49 0.56 0.20 0.47 0.41 27 3 0.11 0.07
dVermin+G 0.47 0.43 37 6 0.16 0.05 0.49 0.44 29 18 0.62 0.07 0.47 0.41 13 3 0.23 0.07
Intra-page bug detection method
iFixdataloss 0.33 0.30 21 8 0.38 0.06 0.44 0.40 52 18 0.35 0.07 0.43 0.39 13 2 0.15 0.04
SetDroid 0.31 0.27 22 8 0.36 0.06 0.44 0.40 98 44 0.45 0.18 0.43 0.39 11 2 0.18 0.04
Genie 0.32 0.31 55 12 0.22 0.09 0.44 0.40 61 22 0.36 0.09 0.43 0.39 27 4 0.15 0.09
Odin 0.33 0.30 57 22 0.39 0.17 0.44 0.40 58 29 0.50 0.12 0.43 0.39 29 6 0.21 0.13
Lint - - 1 1 1.00 0.01 - - 2 1 0.50 0.00 - - 1 1 1.00 0.02
LiveDroid - - 1 1 1.00 0.01 - - 3 1 0.33 0.00 - - 2 0 0.00 0.00
GPT-4 0.39 0.31 55 12 0.22 0.09 0.44 0.40 61 22 0.36 0.09 0.43 0.39 18 2 0.11 0.04
VisionDroid 0.57 0.55 109 55 0.50 0.42 0.65 0.56 177 127 0.72 0.51 0.61 0.59 38 29 0.76 0.64
@@
  • Notes: “Act” is activity coverage, “Cod” is Code coverage. “Rep” is number of bugs reported by tools, “Bug” is true bug in the reported bug, “P” is precision and “R” is the recall. “+A” means it with APE, “+F” means it with Fastbot, “+H” means it with Hunmanoid, and “+G” means it with GPTDroid.
@@
For RQ3, we further evaluate the usefulness of VisionDroid in detecting unseen bugs. We randomly selected 105 popular apps based on app type and download number. We use the same configurations as the previous experiments. Once a bug is spotted, we create an issue report by describing the bug, and report them to the app development team through the issue reporting system or email.
IV-B Baselines
To demonstrate the advantage of VisionDroid, we compare it with 10 common-used and state-of-the-art non-crash bug detection techniques and 4 common-used automated GUI testing techniques. We classify these baselines based on the categories of bugs, including intra-page bug detection approaches and inter-page bug detection approaches.

### v2

@@ II-B Process-aware Monitor Agent
II-B1 Function-based Testing History Construction
@@
The text view depicts the higher-level catalog of testing progress, organized according to the functionalities. A catalog of tested functionalities {Functionality list} with their {Visit number} (visit counts) and {Functionality status} (e.g., fully tested, partially tested). This assists the MLLM in avoiding redundant exploration and identifying untouched functionalities. During testing, the MLLM infers the current functionality being tested through a functionality inquiry prompt, enabling the collection of functionality-specific statistics.
The image view offers a detailed snapshot of the testing records through screenshots, providing a granular view of the testing process. Each GUI page is marked with widgets or actions performed, and screenshots are associated with page descriptions. This ensures alignment between visual and textual information and supports thorough exploration and bug detection.
Figure 6: Example of monitor prompt
II-B2 Monitor Prompt
Fig. 6 demonstrates the monitor prompt. It inputs the testing recording of the current function and the whole testing history, in the form of both text and image, and queries the MLLM about the GUI page description, currently tested function and its status.
II-C Function-aware Detector Agent
The Detector Agent is the vigilant observer of Trident that identifies potential functional bugs by analyzing the logical transitions and interactions within the GUI pages. Upon activation by the Monitor Agent, the Detector Agent first examines the execution sequence of functionality to identify logical inconsistencies for bug detection, in which we design the functionality-driven Chain-of-Thought strategy and enrich the prompt with examples to enhance the performance. The Detector then analyzes the sequence to uncover any potential bug-triggering operations that have not yet been explored. It also identifies any unusual or abnormal operations within the execution sequence that might influence the exploration performance. These results are fed back to the Explorer Agent, which will integrate these insights into the follow-up exploration. This iterative cycle ensures that the GUI testing is thorough and efficient, enhancing the likelihood of uncovering non-crash functional bugs while systematically addressing any anomalies detected during the process.
Figure 7: Example of detector prompt
II-C1 Functionality-driven Chain of Thought (COT)
We design a functionality-driven Chain of Thought (COT) method, which guides the LLM through a structured, sequential analysis of the testing functionality. This ensures that each step is carefully evaluated, reducing the likelihood of hallucination and enhancing the overall effectiveness of the bug detection process.
The MLLM first analyzes the functionality being tested, the expected path, and the actual path taken during testing. This involves the following steps: Functionality Identification: The model identifies the functionality being tested, including its name and description. This helps the model understand the specific functionality it needs to analyze. Expected Path: The model predicts the exploration steps for the functionality based on its description. This step sets a benchmark for the expected behavior. Actual Path: The model reviews the actual sequence of actions and states encountered during testing. Each page and corresponding action are listed sequentially, allowing the model to compare this path against the expected behavior. The prompt also includes an annotated image of the test sequence, with each screenshot labeled according to the type of action (e.g., click, input). This visual representation helps the model correlate the text information with the GUI elements.
II-C2 Enriching Detector Prompt with Examples
@@
It is usually difficult for MLLM to perform as well on domain-specific tasks as ours, and common practice would be employing the in-context learning schema to boost the performance. It provides the MLLM with examples to demonstrate what the instruction is, which enables the MLLM to better understand the task. Following the schema, along with the prompt for the detector, as described in Fig. 7, we additionally provide the MLLM with examples of the non-crash logical bugs.
@@
To achieve this, we first build a basic example dataset of non-crash functional bugs from the issue reports of open-source mobile apps and continuously enlarge it. For each data instance, it records the bug description, bug screenshot and natural language described bug reproduction path which facilitates the MLLM understanding of what the non-crash functional bugs are. We select 200 bugs with the highest quality and complete information as the example data. We also enrich the example dataset with the newly emerged non-crash functional bugs that run on various apps.
@@
Although examples can provide intuitive guidance to the LLM in accomplishing a task, excessive examples might mislead the LLM and cause the performance to decline. Therefore, we design a retrieval-based example selection method to choose the most suitable examples (i.e., the bug page’s description is most similar to the currently tested GUI page’s description) for MLLM. The similarity comparison is based on the activity name of the tested page. Similar to previous works [21], We use Word2Vec [22] to encode the context information of each activity into a sentence embedding, and calculate the cosine similarity between the activity and each data instance in the example dataset. We choose the top-K data instance with the highest similarity score and set K as 5.
II-C3 Detector Prompt
Fig. 7 demonstrates the detector prompt. It consists of the bug examples (as shown in Section II-C2), functionality-driven COT (as shown in Section II-C1) analyzing the functionality being tested, and the real exploration sequence in the form of screenshots. It queries the MLLM about the bug detection, possible bug path, and exception path.
II-D Implementation
@@
In our implementation, we utilize the GPT-4 Vision model [^2] https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4 (gpt-4-vision-preview) to handle both textual and visual information during the testing process. It obtains the view hierarchy file of the current GUI page through UIAutomator [19] to extract text information of the input widgets. Trident is implemented as a fully automated GUI testing tool, which uses or extends the following tools: VirtualBox and the Python pyvbox [23] for running and Android Debug Bridge (ADB) [24] for interacting with the app under test.
To manage the token limit constraints of the GPT-4V API, we combine multiple screenshots from the testing sequences into a single large image. This approach allows us to input comprehensive visual information cohesively, ensuring the model processes all necessary contexts together. On average, each API call consumes approximately 2,000 to 3,000 tokens. We maintain concise representations of the testing history by using functionality names and summaries, reducing the token count. Additionally, when the limit is exceeded, we delete the earlier test records until it satisfies the constraints.
To enhance efficiency and reduce costs, we merge the prompts for Explorer and Monitor into a single request. This integrated prompt simultaneously queries the MLLM about the current exploration state (for explorer) and the tested functionalities (for monitor), which makes the implementation more efficient and cost-effective.
III Experiment Design
  • RQ1: (Coverage and bugs detection performance) How effective of Trident in testing coverage and bug detection?
@@ III Experiment Design
For RQ1, we first present some general views of Trident for coverage of the explorer and bug detection of the detector. Then we compare Trident with commonly-used and state-of-the-art baseline approaches.
@@
  • RQ2: (Ablation Study) What is the contribution of the (sub-) modules of Trident for coverage and bug detection performance?
@@
  • RQ3: (Usefulness Evaluation) How does our proposed Trident work in real-world situations?
For RQ3, we use Trident to automatically explore the app, and detect unseen non-crash functional bugs, and issue the detected bugs to the development team.
III-A Datasets
@@
Meanwhile, we notice that these apps are updated between 2018 and 2021, and some of them are now incompatible with the latest systems. To further investigate the effectiveness of Trident on more recently updated apps and new types of bugs (the baseline dataset only with 5 categories such as “Content related issues”, “Structure related issues”, “Incorrect interaction logic”, etc.), we collect a second dataset (Github dataset) following similar procedures as the baseline dataset, which contain a total of 9 categories of bugs1, such as “Page loading failure”, “Feedback missing”, “Numerical calculation errors”, “Data operation failures”, “Navigation and link errors”, .etc. In detail, we crawl the 50 most popular apps of each category from Google Play [27], and we keep the ones with at least one update after June. 2024, resulting in 437 apps in 13 Google Play categories [27]. Then we manually review each issue report and app, and filter them according to the following criteria: (1) the app wouldn’t constantly crash on the emulator; (2) it can run all baselines; (3) UIAutomator [19] can obtain the view hierarchy file and its GUI screenshot; (4) the issue report involve non-crash bug and it can be manually reproduced for validation; (5) the app is not used in the bug example dataset. Please note that we follow the name of the app to ensure that there is no overlap between the datasets. Finally, 83 apps with 387 non-crash bugs remain for further experiments.
To avoid potential data leakage issues caused by the use of MLLM and further analyze the bug detection capability of Trident, we construct a third dataset (Injection dataset) through bug injection, which can ensure that MLLM has not learned about these bugs in the open-source repositories. Specifically, we obtain the source code of the second dataset apps from GitHub and manually verify whether the app could conduce the bug injection by modifying the code through repackaging. Among them, 33 apps can run normally after bug injection. We conduct the injection based on the bug description, bug repair code, and bug root cause from GitHub, and implement it following the existing bug injection approach [28, 29, 30]. Finally, 33 apps with 72 non-crash bugs remain for further experiments.
The three datasets totally have 590 non-crash bugs for comprehensive evaluation of the effectiveness of Trident.
@@
OwlEyes+M 0.21 0.18 77 10 0.13 0.08 0.33 0.28 140 62 0.44 0.16 0.35 0.31 96 13 0.14 0.18
NightHawk+M 0.21 0.18 63 13 0.21 0.10 0.33 0.28 136 62 0.46 0.16 0.35 0.31 60 15 0.25 0.21
DiffDroid+M 0.21 0.18 61 15 0.25 0.11 0.33 0.28 133 55 0.41 0.14 0.35 0.31 46 12 0.26 0.17
dVermin+M 0.21 0.18 53 4 0.08 0.03 0.33 0.28 98 28 0.29 0.07 0.35 0.31 33 11 0.33 0.15
OwlEyes+A 0.29 0.27 92 18 0.20 0.14 0.43 0.37 161 85 0.53 0.22 0.42 0.4 93 10 0.11 0.14
NightHawk+A 0.29 0.27 86 21 0.24 0.16 0.43 0.37 157 91 0.58 0.24 0.42 0.4 67 16 0.24 0.22
DiffDroid+A 0.29 0.27 87 20 0.23 0.15 0.43 0.37 147 68 0.46 0.18 0.42 0.4 44 7 0.16 0.10
dVermin+A 0.29 0.27 67 6 0.09 0.05 0.43 0.37 101 31 0.31 0.08 0.42 0.4 35 9 0.26 0.13
OwlEyes+F 0.39 0.37 107 17 0.16 0.13 0.47 0.42 180 79 0.44 0.20 0.43 0.39 94 12 0.13 0.17
NightHawk+F 0.39 0.37 100 20 0.20 0.15 0.47 0.42 161 92 0.57 0.24 0.43 0.39 59 15 0.25 0.21
DiffDroid+F 0.39 0.37 94 20 0.21 0.15 0.47 0.42 131 72 0.55 0.19 0.43 0.39 53 12 0.23 0.17
dVermin+F 0.39 0.37 41 8 0.20 0.06 0.47 0.42 90 35 0.39 0.09 0.43 0.39 67 8 0.12 0.11
OwlEyes+H 0.38 0.35 97 12 0.12 0.09 0.48 0.44 157 79 0.50 0.20 0.43 0.39 93 14 0.15 0.19
NightHawk+H 0.38 0.35 88 16 0.18 0.12 0.48 0.44 146 79 0.54 0.20 0.43 0.39 88 13 0.15 0.18
DiffDroid+H 0.38 0.35 75 17 0.23 0.13 0.48 0.44 143 57 0.40 0.15 0.43 0.39 57 15 0.26 0.21
dVermin+H 0.38 0.35 33 5 0.15 0.04 0.48 0.44 82 40 0.49 0.10 0.43 0.39 41 8 0.20 0.11
OwlEyes+G 0.47 0.43 99 17 0.17 0.13 0.51 0.47 162 79 0.49 0.20 0.47 0.41 95 14 0.15 0.19
NightHawk+G 0.47 0.43 95 18 0.19 0.14 0.51 0.47 157 96 0.61 0.25 0.47 0.41 68 12 0.18 0.17
DiffDroid+G 0.47 0.43 85 16 0.19 0.12 0.51 0.47 142 76 0.54 0.20 0.47 0.41 47 9 0.19 0.13
dVermin+G 0.47 0.43 37 6 0.16 0.05 0.51 0.47 75 47 0.63 0.12 0.47 0.41 42 9 0.21 0.13
iFixdataloss 0.33 0.3 21 8 0.38 0.06 0.48 0.43 101 45 0.45 0.12 0.43 0.39 41 10 0.24 0.14
SetDroid 0.31 0.27 22 8 0.36 0.06 0.48 0.43 164 64 0.39 0.17 0.43 0.39 47 11 0.23 0.15
Genie 0.32 0.31 55 12 0.22 0.09 0.48 0.43 123 46 0.37 0.12 0.43 0.39 43 8 0.19 0.11
Odin 0.33 0.3 57 22 0.39 0.17 0.48 0.43 107 59 0.55 0.15 0.43 0.39 57 11 0.19 0.15
Lint - - 1 1 1.00 0.01 - - 54 30 0.56 0.08 - - 2 2 1.00 0.03
LiveDroid - - 1 1 1.00 0.01 - - 58 24 0.41 0.06 - - 1 1 1.00 0.01
GPT-4 0.39 0.31 55 12 0.22 0.09 0.47 0.45 123 48 0.39 0.12 0.43 0.39 35 8 0.23 0.11
AppAgent 0.41 0.38 47 6 0.13 0.05 0.44 0.43 113 12 0.11 0.03 0.39 0.37 31 4 0.13 0.06
Trident 0.57 0.55 109 55 0.50 0.42 0.65 0.62 278 201 0.72 0.52 0.61 0.59 67 47 0.70 0.65
@@
  • Notes: “Act” is activity coverage, “Cod” is Code coverage. “Rep” is number of bugs reported by tools, “Bug” is true bug in the reported bug, “P” is precision and “R” is the recall. “+M” means it with Monkey, “+A” means it with APE, “+F” means it with Fastbot, “+H” means it with Hunmanoid, and “+G” means it with GPTDroid.
@@
For RQ3, we further evaluate the usefulness of Trident in detecting unseen bugs. We randomly selected 187 popular apps based on app type and download number. We use the same configurations as the previous experiments. Once a bug is spotted, we create an issue report by describing the bug, and report them to the app development team through the issue reporting system or email.
III-B Baselines
To demonstrate the advantage of Trident, we compare it with 12 common-used and state-of-the-art GUI testing techniques.